5.92(g) of hydrated Feso4 when heated to a constant mass gave 3.24(g) of anhydrous salt. Calculate the number if moles off water crystallization in the hydrated salt (H=1, O=16, S=32, Fe=52)

1
Expert's answer
2022-03-17T14:58:09-0400

Solution:

hydrated FeSO4 = FeSO4⋅nH2O

anhydrous salt = FeSO4


The molar mass of anhydrous FeSO4 is (56 + 32 + 16×4) = 152 g/mol

Therefore,

Moles of FeSO4 = (3.24 g FeSO4) × (1 mol FeSO4 / 152 g FeSO4) = 0.0213 mol FeSO4


Schematic chemical equation:

FeSO4⋅nH2O ⟶ FeSO4 + nH2O

According to the principle of mass conservation:

Mass of FeSO4⋅nH2O = Mass of FeSO4 + Mass of H2O

Therefore,

Mass of H2O = Mass of FeSO4⋅nH2O − Mass of FeSO4 = 5.92 g − 3.24 g = 2.68 g


The molar mass of water (H2O) is (1×2 + 16) = 18 g/mol

Therefore,

Moles of H2O = (2.68 g H2O) × (1 mol H2O / 18 g H2O) = 0.1489 mol H2O


The mole ratio between FeSO4 and H2O in the hydrated salt:

Moles of FeSO4 : Moles of H2O = 0.0213 : 0.1489= 1 : 6.99 = 1 : 7

Thus, the formula of the hydrated salt is FeSO4⋅7H2O

The number of moles of water crystallization in the hydrated salt is '7'


Answer: The number of moles of water crystallization in the hydrated salt is '7'